OpenStudy (anonymous):

What is the slope of the line that passes through the given points? (6,-1) and (-3,-1) 5 minutes ago

jimthompson5910 (jim_thompson5910):

use the slope formula m = (y2-y1)/(x2-x1) m = (-1-(-1))/(-3-6) m = (-1+1)/(-3-6) m = ???

OpenStudy (anonymous):

Both y values are -1, and therefore there can be no slope/gradient along these points so basically, slope = 0

OpenStudy (anonymous):

So it's 0 not undefined

jimthompson5910 (jim_thompson5910):

yes the slope is 0 this line is a horizontal line

OpenStudy (anonymous):

is that not what I just said?

OpenStudy (anonymous):

Yeah but I was askin you don't have to be rude about it